How do you solve 2x-4y=7 to put it in slope-intercept form? I tried this:subracting 2x on both sides, then writing it out as -4y-2x=7. Is that correct?
Answer by rfer(16322)

2x-4y=7
subtract 2x from both sides
2x-2x-4y=-2x+7
-4y=-2x+7
multyiply through by -
4y=2x-7
-------------------
Does that help getting you going?